OpenStudy (m3ga-bit3):

What is (2+y)³ written out?

OpenStudy (m3ga-bit3):

@marlingdfolkhayes

OpenStudy (m3ga-bit3):

@ganeshie8

OpenStudy (m3ga-bit3):

@protected

OpenStudy (m3ga-bit3):

A. 8 + 12y + 6y² + y³ B. 6 + 8y + 4y² + y³

OpenStudy (anonymous):

Hang on, I gotcha.

ganeshie8 (ganeshie8):

\[(2+y)^3 = (2+y)(2+y)(2+y)\]

OpenStudy (m3ga-bit3):

yes

OpenStudy (alexandervonhumboldt2):

2^3+3*2^2*y+3*2*y^2+y^3

OpenStudy (alexandervonhumboldt2):

simplify now

OpenStudy (alexandervonhumboldt2):

(a+b)^3=a^3+3a^2b+3ab^2+b^3

OpenStudy (alexandervonhumboldt2):

did you got it?

OpenStudy (alexandervonhumboldt2):

8+12y+6y^2+y^3 is the answer

OpenStudy (anonymous):

Eh, i left and they beat me to it. You should reward @AlexandervonHumboldt2 :)

OpenStudy (alexandervonhumboldt2):

ME?

OpenStudy (anonymous):

You answered it didnt you lol.

OpenStudy (alexandervonhumboldt2):

thx

OpenStudy (m3ga-bit3):

nice job guys i appreaciate it